\int \:\frac{x+sinx}{1+cosx}dx
1
Expert's answer
2020-10-11T18:04:26-0400
"I=\int \:\frac{x+\sin x}{1+\cos x}dx={\displaystyle\int}\left(\dfrac{\sin\left(x\right)}{\cos\left(x\right)+1}+\dfrac{x}{\cos\left(x\right)+1}\right)\mathrm{d}x"

Thus,

Applying linearity we get,


"I={\displaystyle\int}\dfrac{\sin\left(x\right)}{\cos\left(x\right)+1}\,\mathrm{d}x+{\displaystyle\int}\dfrac{x}{\cos\left(x\right)+1}\,\mathrm{d}x"

Now, consider


"I_1={\displaystyle\int}\dfrac{\sin\left(x\right)}{\cos\left(x\right)+1}\,\mathrm{d}x"

Let,


"u=1+\cos x\implies \sin x dx =-du"

, hence


"I_1=-{\displaystyle\int}\dfrac{1}{u}\,\mathrm{d}u=-\ln u=-\ln (1+\cos x)+c_1"

Consider,


"I_2={\displaystyle\int}\dfrac{x}{\cos\left(x\right)+1}\,\mathrm{d}x"

Thus, after rewriting the above integrand we get


"I_2={\displaystyle\int}\dfrac{x\csc\left(x\right)}{\csc\left(x\right)+\cot\left(x\right)}\,\mathrm{d}x"

Now, integrate by parts, take "f=x" , "g'=\dfrac{\csc\left(x\right)}{\csc\left(x\right)+\cot\left(x\right)}"

And apply

"{\int}\mathtt{f}\mathtt{g}' = \mathtt{f}\mathtt{g} - {\int}\mathtt{f}'\mathtt{g}"

Thus, "f'=1" and


"g={\displaystyle\int}\dfrac{\csc\left(x\right)}{\csc\left(x\right)+\cot\left(x\right)}\,\mathrm{d}x\\"

Now, substitute "u=\dfrac{1}{\csc\left(x\right)+\cot\left(x\right)}" ,hence "\mathrm{d}x=-\dfrac{\left(\csc\left(x\right)+\cot\left(x\right)\right)^2}{-\csc^2\left(x\right)-\cot\left(x\right)\csc\left(x\right)}\,\mathrm{d}u"

Therefore,


"g=\int du=u\implies g=\dfrac{1}{\csc\left(x\right)+\cot\left(x\right)}"

Hence,


"I_2=\dfrac{x}{\csc\left(x\right)+\cot\left(x\right)}-{\displaystyle\int}\dfrac{1}{\csc\left(x\right)+\cot\left(x\right)}\,\mathrm{d}x"

Now, consider


"I_3={\displaystyle\int}\dfrac{1}{\csc\left(x\right)+\cot\left(x\right)}\,\mathrm{d}x"

After simplification of "I_3" we get


"I_3=-{\displaystyle\int}-\dfrac{1}{\cos\left(x\right)+1}{\sin x} dx"

Take

"v=1+\cos x\implies -\sin xdx=dv"

Hence,


"I_3=-\int \frac{1}{v}dv=-\ln v=-\ln(1+\cos x)+c_2"



Thus,

"I_2=\dfrac{x}{\csc\left(x\right)+\cot\left(x\right)}-I_3\\
\implies I_2=\dfrac{x}{\csc\left(x\right)+\cot\left(x\right)}+\ln (1+\cos x)+c_2"

Therefore,


"I=I_1+I_2\implies I=\dfrac{x}{\csc\left(x\right)+\cot\left(x\right)}+c"

Where, constant "c=c_1+c_2"